OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity a to the 2nd power over b to the 3rd power all raised to the 4th power. a to the 6th power over b to the 7th power b12 a8b15 the quantity a to the 8th power over b to the 12th power

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

is it d?\[\frac{ a^8 }{ b^12 }\]

OpenStudy (anonymous):

b^12, so like an exponent of 12, not 1...

OpenStudy (jdoe0001):

hmm what makes you think is d anyway?

OpenStudy (anonymous):

idk, I originally thought it was a, but idk

OpenStudy (jdoe0001):

\(\bf (a^n)^m\implies a^{n\cdot m}\quad thus \\ \quad \\ \left( \cfrac{a^2}{b^3} \right)^{\color{brown}{ 4}}\implies \left( \cfrac{a^{2\cdot {\color{brown}{ 4}}}}{b^{3\cdot {\color{brown}{ 4}}}} \right)\)

OpenStudy (jdoe0001):

so...should be obvous I'd think